# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v2

@@ VI Discussion
The FHN model’s dynamics are influenced by a range of factors including external driving, internal feedback, inter-system coupling, noise, and time delays. These elements are important not only in the model’s inherent complexity but also in its extensive applicability across physical and biological domains. While in our review, we have studied in some detail the dynamics of single and coupled FHN systems, it has not extensively covered the roles of external driving, self-feedback, time delays, and noise, each of which merits further exploration for a comprehensive understanding of the FHN model’s dynamics and potential. Additionally, the integration of the FHN model within multiphysics frameworks, which encapsulate diverse physical interactions, has not been explored. Such integrative models are particularly important in the complex simulation of biological phenomena, with cardiac and neuronal examples highlighting the FHN model’s adaptability and significance in these areas.
External Driving.: Introducing an external stimulus to the FHN system can induce resonance or forced oscillations, when the stimulus frequency aligns with the system’s inherent frequency. This aspect is pivotal in understanding biological rhythms and has applications in electronic and mechanical systems, as well as in neuroscience for modeling brain rhythms and treating disorders like Parkinson’s. Researchers have looked into the dynamics of externally driven FHN models, revealing their potential for controlling excitable behaviors Ullner et al. (2003); Pankratova et al. (2005); Kuramoto et al. (1998). Of course, the external driving could also come from another FHN system, leading to two autonomous dynamical systems in a master–slave configuration. Synchronisation has been studied in such unidirectionally coupled models, also in the presence of additional external aperiodic forcing Ciszak et al. (2003).
@@
Counterintuitively, random perturbations, especially when combined with weak deterministic stimuli, can introduce order in temporal and spatial domains of nonlinear systems, a phenomenon known as stochastic resonance. This effect, where noise enhances system response to external perturbations, underscores noise’s constructive potential in nonlinear dynamics. Coherence resonance occurs when an excitable system exhibits maximal signal regularity at optimal noise levels, even without external driving forces. This phenomenon, alongside stochastic resonance, has been pivotal in understanding dynamics near bifurcation points and within bistable and oscillatory systems. Also in the context of the FHN model, it has sparked significant interest for its implications in synchronization Yu et al. (2021); Neiman et al. (1999); Brandstetter et al. (2010); Makarov et al. (2001), resonance Lindner and Schimansky-Geier (1999); Toral et al. (2003a); Lindner (2004); Wiesenfeld et al. (1994); Longtin (1993); Perc (2005b); Brandstetter et al. (2010); Semenova et al. (2016); Zaks et al. (2005); Pankratova et al. (2005), and the formation of complex spatio-temporal patterns like spirals Perc (2007); Toral et al. (2003a); García-Ojalvo and Schimansky-Geier (1999).

### v3

@@ VI Discussion
The dynamics of the FHN model are influenced by various factors, including external driving, internal feedback, inter-system coupling, noise, and time delays. These elements contribute not only to the model’s inherent complexity but also to its extensive applicability across physical and biological domains. While our review has examined in detail the dynamics of single and coupled FHN systems, it has not extensively covered the roles of external driving, self-feedback, time delays, and noise. Each of these aspects warrants further exploration for a comprehensive understanding of the FHN model’s dynamics and potential. Living systems are typically non-equilibrium, active, and multi-scale complex systems that often operate in transient states away from stable attractors. The FHN model is an excellent candidate to serve as a testbed for developing new methods to analyze such transient dynamics and non-reciprocal interactions. Integrating the FHN model within multiphysics frameworks, which encapsulate diverse physical interactions, is essential for advancing our understanding of these systems. Such new techniques and integrative models are particularly important in the complex simulation of biological phenomena. Cardiac and neuronal examples, in particular, highlight the FHN model’s adaptability and significance in these areas.
External Driving.: Introducing an external stimulus to the FHN system can induce resonance or forced oscillations, when the stimulus frequency aligns with the system’s inherent frequency. This aspect is pivotal in understanding biological rhythms and has applications in electronic and mechanical systems, as well as in neuroscience for modeling brain rhythms and treating disorders like Parkinson’s. Researchers have looked into the dynamics of externally driven FHN models, revealing their potential for controlling excitable behaviors Ullner et al. (2003a); Pankratova et al. (2005); Kuramoto et al. (1998). Of course, the external driving could also come from another FHN system, leading to two autonomous dynamical systems in a master–slave configuration. Synchronisation has been studied in such unidirectionally coupled models, also in the presence of additional external aperiodic forcing Ciszak et al. (2003).
@@
Counterintuitively, random perturbations, especially when combined with weak deterministic stimuli, can introduce order in temporal and spatial domains of nonlinear systems, a phenomenon known as stochastic resonance. This effect, where noise enhances system response to external perturbations, underscores noise’s constructive potential in nonlinear dynamics. Coherence resonance occurs when an excitable system exhibits maximal signal regularity at optimal noise levels, even without external driving forces. This phenomenon, alongside stochastic resonance, has been pivotal in understanding dynamics near bifurcation points and within bistable and oscillatory systems. Also in the context of the FHN model, it has sparked significant interest for its implications in synchronization Yu et al. (2021); Neiman et al. (1999); Brandstetter et al. (2010); Makarov et al. (2001), resonance Lindner and Schimansky-Geier (1999); Toral et al. (2003a); Lindner (2004); Wiesenfeld et al. (1994); Longtin (1993); Perc (2005b); Brandstetter et al. (2010); Semenova et al. (2016); Zaks et al. (2005); Pankratova et al. (2005); Sagués et al. (2007), the formation and stability of complex spatio-temporal patterns like spirals Perc (2007); Toral et al. (2003a); García-Ojalvo and Schimansky-Geier (1999); Sagués et al. (2007); Ullner et al. (2003b); García-Ojalvo et al. (2001).
Transient dynamics. Many techniques developed in nonlinear dynamics focus on analyzing the stability and structure of attracting solutions. However, living systems, in particular, are most often in transient Hastings et al. (2018); Rabinovich et al. (2008); Verd et al. (2014) and continuously integrate time-varying, noisy information from both internal and external sources. Currently, there is a lack of a clear framework to characterize such long transient dynamics of system state trajectories. The need to adapt the mathematical formalism when describing biological systems in changing environments has recently been emphasized Koch and Koseska (2024); Koch et al. (2024); Nandan and Koseska (2023).
Nonreciprocally coupled systems. Non-reciprocity is a common feature out of equilibrium, where effective interactions between agents violate Newton’s third law. Such interactions are typically observed in living systems, which are often active, with interacting agents converting free energy into directed movement. Despite the widespread importance of non-reciprocal interactions, a general framework for describing their effects is still lacking, hindering our ability to control and exploit them. In recent years, non-reciprocally coupled systems have received growing attention Fruchart et al. (2021); Dinelli et al. (2023); Saha et al. (2020); Brauns et al. (2020); Brauns and Marchetti (2024); Frohoff-Hülsmann and Thiele (2023). Studies on particle and spin models, as well as effective continuum descriptions, have demonstrated that non-reciprocal systems generically give rise to spontaneous currents and non-equilibrium patterns that are not typically seen in the absence of such interactions. These phenomena occur in a wide range of pattern-forming systems, including mass-conserving reaction-diffusion systems. Such systems can be mapped to a mass-conserving FHN model, similar to the non-reciprocal Cahn-Hilliard model, which exhibits rich spatio-temporal behaviors in oscillatory and excitable media Saha et al. (2020); Brauns et al. (2020); Brauns and Marchetti (2024); Frohoff-Hülsmann and Thiele (2023).